### v1

# CaT-GNN: Enhancing Credit Card Fraud Detection via Causal Temporal Graph Neural Networks

Abstract
Credit card fraud poses a significant threat to the economy. While Graph Neural Network (GNN)-based fraud detection methods perform well, they often overlook the causal effect of a node’s local structure on predictions. This paper introduces a novel method for credit card fraud detection, the Causal Temporal Graph Neural Network (CaT-GNN), which leverages causal invariant learning to reveal inherent correlations within transaction data. By decomposing the problem into discovery and intervention phases, CaT-GNN identifies causal nodes within the transaction graph and applies a causal mixup strategy to enhance the model’s robustness and interpretability. CaT-GNN consists of two key components: Causal-Inspector and Causal-Intervener. The Causal-Inspector utilizes attention weights in the temporal attention mechanism to identify causal and environment nodes without introducing additional parameters. Subsequently, the Causal-Intervener performs a causal mixup enhancement on environment nodes based on the set of nodes. Evaluated on three datasets, including a private financial dataset and two public datasets, CaT-GNN demonstrates superior performance over existing state-of-the-art methods. Our findings highlight the potential of integrating causal reasoning with graph neural networks to improve fraud detection capabilities in financial transactions.
1 Introduction
The substantial damages wrought by financial fraud continue to garner ongoing focus from academic circles, the business sector, and regulatory bodies Jiang et al. (2016); Aleksiejuk and Hołyst (2001). Fraudsters masquerade as ordinary users and attack transactions made with credit cards Ileberi et al. (2022), which may inflict substantial economic losses and pose a severe threat to sustainable economic growth AlFalahi and Nobanee (2019). Consequently, effective detection of financial fraud is imperative for safeguarding the economy and consumer security.
In the financial deception realm, identifying credit card fraud has garnered considerable attention among both industry and academia Bhattacharyya et al. (2011). Traditional approaches to detecting fraudulent activities typically entail meticulous examination of each transaction for irregularities, employing predefined criteria such as verification against lists of compromised cards or adherence to established transaction thresholds Maes et al. (2002); Fu et al. (2016). However, the aforementioned anti-fraud systems, based on expert prior and rules, are often susceptible to exploitation by fraudsters, who can circumvent detection by crafting ingenious transaction methods that elude the system’s scrutiny of illicit activities. Toward this end, predictive modeling has been introduced, aiming to autonomously identify patterns that suggest fraudulent activity and calculate a corresponding risk score.
Figure 1: The model overview. First Stage (discovery): we utilize an attention map in the attention temporal network to identify causal nodes and environment nodes. Second Stage: Intervention, we apply causal mix-up enhancement to the environment nodes.
Currently, state-of-the-art predictive models are focused on using deep learning methods, capturing potential illegal patterns in a data-driven manner Fu et al. (2016); Dou et al. (2020). For instance, Liu et al. (2021) introduces PC-GNN, a Graph Neural Network approach that effectively handles class imbalance in graph-based fraud detection by selectively sampling nodes and edges, particularly focusing on the minority class. Moreover, Xiang et al. (2023) leverages transaction records to construct a temporal transaction graph, applying a Gated Temporal Attention Network to effectively learn transaction representations and model fraud patterns. Unfortunately, i) these methods often overlook the intrinsic patterns and connections within the data due to a lack of consideration for local structure consistency; ii) they lack the ability to uncover the causal nature of each specific case, which leads to inadequate differentiation between the attributes of causal nodes and environment nodes, thereby impairing the model’s generalization capabilities; iii) they lack interpretability in making specific predictions.
In this paper, we introduce a novel Causal Temporal Graph Neural Network, termed CaT-GNN, aiming at providing an interpretable paradigm for credit card fraud detection. Guided by the currently popular causal invariant learning techniques Chang et al. (2020); Liu et al. (2022), CaT-GNN’s primary objective is to unveil the inherent correlations in the transaction attribute data of nodes within available temporal transaction graphs, thereby offering interpretability for complex transaction fraud problems.
To unravel causal correlations, specifically, we decompose the algorithmic process of CAT-GNN into two stages - discovery and intervention. The goal of the discovery stage is to identify potential causal components within observed temporal graph data, where we introduce a causal temporal graph neural network for modeling. Utilizing the popular node-attention metrics Veličković et al. (2017); Xiang et al. (2023), we employ attention score to locate key nodes, designated as causal and environment nodes. In the intervention process, we aim to reasonably enhance potential environment nodes. This approach is designed to align with and perceive the underlying distribution characteristics in explicit fraud networks, thereby boosting our temporal GNN’s ability to identify and understand problematic nodes. Furthermore, drawing inspiration from Wang et al. (2020), to ensure that causal interventions between nodes do not interfere with each other, we create parallel universes for each node. Consequently, the model is exposed to a wider potential data distribution, providing insights for fraud prediction with a causal perspective. This process can further be understood as a back-door adjustment in causal theory Pearl (2009); Pearl and Mackenzie (2018). The contributions of this paper are summarized as follows:
  • We propose a novel method, CaT-GNN, that embodies both causality and resilience for the modeling of credit card fraud detection. By harnessing causal theory, known for its interpretability, CaT-GNN enables the model to encompass a wider potential data distribution, thereby ensuring its exceptional performance in this task.
  • CaT-GNN, characterized by its refined simplicity, initially identifies causal nodes and subsequently refines the model into a causally coherent structure. It aims to achieve invariance in attribute information and temporal features through semi-supervised learning, thereby providing a bespoke and robust foundation for targeted tasks.
  • We evaluate CaT-GNN on three representative datasets, including a private financial benchmark, and the other two are public settings. Extensive experiments show that our proposed method outperforms the compared state-of-the-art baselines in credit card fraud detection, thanks to the casual intervention of the node causal augment.
2 Preliminaries
Definition 1. (Multi-Relation Graph)
The Multi-Relation Financial Graph $\mathcal{G}$ is defined as $\mathcal{G}=(\mathcal{V,E,X,Y})$, where $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{N}\right\}$ represents the set of nodes, with $N=|\mathcal{V}|$ indicates the total number of nodes. $\mathcal{X}\in\mathbb{R}^{N\times d}$ denotes the node features with $x_{i}\in\mathbb{R}^{d}$ as its entry for node $v_{i}$, $d$ is the feature dimension. Each node $v_{i}$ is assigned a label $y_{i}\in\mathcal{Y}$, which is a binary variable with the value in $\left\{0,1\right\}$. =$\mathcal{E}=\left\{\mathcal{E}_{1},\cdots,\mathcal{E}_{R}\right\}$ signifies the set of edges, partitioned into $R$ distinct types of relations.
Definition 2. (Graph-based Fraud Detection)
The graph-based fraud detection problem is defined on the multi-relation graph $\mathcal{G}=(\mathcal{V,E,X,Y)}$. For such a problem, each node $v_{i}$ represents the target entity such as a transaction record, and has a label $y_{i}\in Y$, where $y_{i}=0$ represents benign and $y_{i}=1$ represents fraud. The objective of graph-based fraud detection is to identify fraud nodes that stand out distinctly from the non-fraudulent, or benign, nodes within a multi-relational graph $\mathcal{G}$. This task is effectively approached as a binary classification problem focused on nodes within the graph $\mathcal{G}$. And the surprised Binary Cross-Entropy Loss function is:$\quad\min_{\Theta}\mathcal{L}=-\frac{1}{B}\sum_{u=1}^{B}\left(y_{i}^{\top}\log\left(\hat{y}_{i}\right)+\left(1-y_{i}\right)^{\top}\log\left(1-\hat{y}_{i}\right)\right)+\eta||\Theta||^{2},$ where $B$ is the batch size, $\hat{y}_{i}\in[0,1]$ is the predicted probability, and $y_{i}\in\{0,1\}$ is the ground truth. $\Theta$ is the parameter of the GNN predictor.
3 Methodology
In Section 3.1, we explore the motivation behind our approach, emphasizing the crucial role of understanding the local structure and causal relationships within transaction data to improve detection accuracy. Section 3.2 introduces our two-phase method: discovery and intervention. Section Section 3.2 provides the causal theory support.
3.1 Motivation
Figure 2: Motivation. The original prediction incorrectly identifies a fraudster (central node labeled $x_{i}$) as benign, as does the state-of-the-art GTAN model. Following our causal intervention, the prediction is correctly adjusted to identify $x_{i}$ as a fraudster. Green: benign users, red: fraudsters, gray: unlabeled nodes.
Figure 3: The depiction of the proposed model’s architecture, featuring a causal temporal graph attention mechanism, alongside the theoretical support for backdoor adjustment.
Taking the arXiv Hu et al. (2020) dataset as an example, real-world graphs often exhibit locally variable structures, that is, the distribution of node attributes differs from the distribution of local structural properties Feng et al. (2021). We observe that this phenomenon is also prevalent in the financial sector, where cunning fraudsters may disguise themselves through various means (such as feature camouflage and relationship disguise) to connect with users who have a good credit transaction history Dou et al. (2020). In such scenarios, if we simply aggregate node information and neighbor information together, it is likely to obscure the suspiciousness of the fraudsters, which contradicts our objective. This situation tends to yield poorer training outcomes, especially in a semi-supervised learning environment with limited labeled data. Existing methods do not incorporate causal factors into credit card fraud modeling, resulting in models that fail to learn the intrinsic connections of node attributes. This oversight further leads to the neglect of causal attribute structure differences on test nodes, thereby reducing the model’s generalizability. By comprehensively examining the confounding variables, we are able to significantly alleviate the aforementioned issue, as illustrated in Figure 2. This strategy is the cornerstone of our framework and is also known as the “backdoor adjustment” technique Pearl (2009); Pearl and Mackenzie (2018).
3.2 Discovering & Intervention
Based on the motivation, we adopt a causal perspective to analyze the attribute aggregation process and formalize principles for distinguishing between causal and non-causal elements within local structures. We first introduce the discovery process to effectively examine the causal and environment nodes within the current node’s local structure. In response, we refine the temporal attention graph network mechanismXiang et al. (2022) into a causal temporal GAT mechanism as shown in the upper half of Figure 3. This refinement introduces two key components designed to accurately identify both environmental and causal nodes, which enhances our ability to understand and manipulate the local structural dynamics more effectively.
In the context of temporal transaction graphs, we maintain a set of transaction records, denoted as $R=[r_{t1},r_{t2},\cdots,r_{ti}]$, alongside their embeddings $X=[x_{t1},x_{t2},\cdots,x_{ti}]$ obtained via a projection layer. As demonstrated in Shi et al. (2020), GNNs are capable of concurrently propagating attributes and labels. Consequently, we integrate fraud labels as an embedded feature within the attribute embedding $x_{ti}$, employing masking techniques to prevent label leakage Xiang et al. (2023). However, this aspect does not constitute the primary focus of our research.
Causal-Inspector: We design a Causal-Inspector to identify causal and environment nodes as shown in the bottom left corner of Figure 3. To aggregate information efficiently, we employ the aforementioned causal temporal graph attention mechanism, which allows for dynamic information flow based on the temporal relationships among transactions. Leveraging a multi-head attention mechanism, we compute temporal attention scores that serve as weights for each neighboring node, facilitating the assessment of each neighbor’s causal importance, which can be formulated as follows:
$\alpha_{x_{i},x_{j}}^{h}=\frac{\exp\bigl{(}\text{LeakyReLU}(W_{a}^{T}[x_{i}\oplus x_{j}])\bigl{)}}{\sum_{j\in N(x_{i})}\exp(\text{LeakyReLU}\bigl{(}W_{a}^{T}[x_{i}\oplus x_{j}])\bigl{)}},$ (1)
where $W_{a}$ is a learnable weight matrix, $\alpha_{x_{i},x_{j}}$ represents the attention weight of node $x_{i}$ with respect to node $x_{j}$ in one head, which determines the importance of node $x_{i}$ relative to node $x_{j}$. The $\oplus$ symbol represents the concatenation operation. $N(x_{i})$ is the set of temporal neighboring nodes of node $x_{i}$. In order to quantify the importance of each node $x_{j}$ we aggregate the attention weights $\alpha_{x_{i},x_{j}}^{h}$ from each attention head and compute the average to determine the final weight of the node. Then, based on its final weight, we calculate its normalized importance:
$\bar{\alpha}_{x_{i},x_{j}}=\frac{1}{H}\sum_{h=1}^{H}\alpha_{x_{i},x_{j}}^{h}\rightarrow\hskip 5.69046ptI_{j}=\bar{\alpha}_{x_{i},x_{j}}/\sum_{j=1}^{N(x_{i})}\bar{\alpha}_{x_{i},x_{j}}$ (2)
where $H$ is the total number of attention heads and $I_{i}=[I_{j1},\cdots,I_{jN(x_{i})}]$ represents the set of importance of each node with respect to $x_{i}$. This formula calculates the normalized importance weight $I_{j}$, representing the importance of node $x_{j}$ by compiling the contributions from all attention heads, thus providing a comprehensive measure of node significance. To segregate the nodes into environmental and causal categories, we introduce a proportion parameter $r_{e}$, ranging between 0 and 1, which denotes the fraction of nodes to be earmarked as environment nodes. This approach affords us the flexibility to select environment nodes tailored to the specific exigencies of the graph. We use the $\text{argmin}(\cdot)$ function to select the $\lceil r_{e}N\rceil$ nodes with the lowest importance scores as environment nodes. Therefore, a ranking function $R$ is defined to map $I_{j}$ to its rank among all node importance scores. Then, we determine the environment set $S_{e}$ as:
${S_{e}}=\bigl{\{}x_{j}|\mathop{\text{argmin}}\limits_{j=1,\cdots,\lfloor r_{e}N\rfloor}R(I_{j})\bigl{\}}.$ (3)
The remaining nodes, those not in ${S_{e}}$, naturally form the set of causal nodes $S_{c}$.This method ensures that nodes with the lowest importance scores are precisely selected as environmenta nodes according to the proportion $r_{e}$, while the rest serve as causal nodes. Due to the differences between test and training distributions Feng et al. (2021), CaT-GNN is dedicated to perceiving the essence of temporal graph data, thereby enhancing generalization capabilities and robustness.
Causal-Intervener: We design a Causal Intervener as shown in the bottom right corner of Figure 3, which employs a transformative mixup strategy known as a causal mixup, that blends environment nodes with a series of causally significant nodes. Given an environmental node $x_{j}\in S_{e}$, We select the causal nodes $\{x_{c1},x_{c2},\cdots,x_{ck}\}$ with the highest importance scores, which are computed as outlined in the Causal-Inspector, from the causal set $S_{c}$ at a proportion of $r_{c}$. The causal mixup is then executed by linearly combining the environmental node with the selected causal nodes, weighted by their respective coefficients $\{a_{j},a_{1},a_{2},\cdots,a_{k}\}$, which are learned through a dedicated linear layer:
$x^{\prime}_{j}=a_{j}x_{j}+\sum_{i=1}^{k}a_{i}x_{ci},$ (4)
where $x^{\prime}_{j}$ is the causally mixed environmental node, $k$ is the number of selected causal nodes, $a_{j}$ is the self-weight of the environmental node reflecting its inherent causal significance, and $a_{i}$ is the causal node weight. These weights are normalized such that $a_{j}+\sum_{i=1}^{k}a_{i}=1$. The incorporation of the causal mixup enhances the robustness of the model against distributional shifts by embedding a richer causal structure within the environmental node. By adapting the causal structure to the environmental context, the Causal-Intervener aims to mitigate the disparity between training and test distributions, thus bolstering the model’s generalizability. Finally, we aggregate the information, and the outputs of multiple attention heads are concatenated to form a more comprehensive representation:
$\mathcal{H}=\sum_{x_{i}\in\mathcal{X}}\sigma\bigl{(}\sum_{x_{j}\in N(x_{i})}\alpha_{x_{i},x_{j}}x_{j}^{\prime}\bigl{)},$ (5)
$\mathcal{M}=\bigl{(}\mathcal{H}_{1}\oplus\cdots\oplus\mathcal{H}_{\text{h}}\bigl{)}W_{c},$
where $W_{c}$ is a learnable weight matrix, $\mathcal{H}$ is a attention head, $\mathcal{M}$ denotes the aggregated embeddings. It is important to highlight that the causal intervention result $x_{j}^{\prime}$ on an environmental node $x_{j}$ with respect to $x_{i}$ is essentially a duplicate of $x_{j}$ and does not modify $x_{j}$ itself. This distinction is crucial as it guarantees that the process of augmenting central nodes within individual local structures remains mutually non-disruptive. By preserving the original state of $x_{j}$, we ensure that enhancements applied to central nodes in one local structure do not adversely affect or interfere with those in another, maintaining the integrity and independence of local structural enhancements Wang et al. (2020).
3.3 Causal Support of CaT-GNN
In elucidating the causal backbone of CaT-GNN, we invoke causal theory to formulate a Structural Causal Model (SCM) as propounded by Pearl (2009). This framework scrutinizes four distinct elements: the inputs node attribute $X$, the truth label $Y$ decided by both the attribute of causal nodes of $X$ symbolized as $\tilde{C}$, and the confounder $E$, emblematic of the attribute of environment nodes. The causal interplay among these variables can be articulated as follows:
  • $\tilde{C}\xleftarrow{}$$\xrightarrow{}$. The local structure of node attribute $X$ is composed of causal nodes attributes $\tilde{C}$ and environment nodes attributes $E$.
  • $\tilde{C}\xrightarrow{}$$\ \xleftarrow{}$. The causal attributes $\tilde{C}$ actually determine the true value $Y$, however, the environmental attributes $E$ also affect the prediction results, causing spurious associations.
Do-calculus Pearl (2009) is a trio of rules within the causal inference framework that facilitates the mathematical deduction of causal effects from observed data. These rules enable manipulation of $\text{do}(\cdot)$ operator expressions, essential for implementing interventions in causal models:
$P\bigl{(}Y|\text{do}(\hat{C}),E\bigl{)}=P(Y|\hat{C}),$ (6)
$P\bigl{(}Y|\text{do}(\hat{C}),\text{do}(E)\bigl{)}=P\bigl{(}Y|\text{do}(\hat{C}),E\bigl{)},$
$P\bigl{(}\hat{C}|\text{do}(Y)\bigl{)}=P\bigl{(}\hat{C}|Y\bigl{)}.$
Typically, a model $M_{\theta}$ that is trained using Empirical Risk Minimization (ERM) may not perform adequately when generalizing to test data distribution $P_{test}$. These shifts in distribution are often a result of changes within environment nodes, necessitating the need to tackle the confounding effects. As illustrated in Figure 3, we apply causal intervention to enhance the model’s generalizability and robustness. To this end, our approach utilizes do-calculus Pearl (2009) on the variable $C$ to negate the influence of the backdoor path $E\rightarrow Y$ by estimating $P\bigl{(}Y|\text{do}(\hat{C})\bigl{)}=P_{m}(Y|\hat{C})$:
$P\bigl{(}Y|\text{do}(\hat{C})\bigl{)}$ $=\sum_{i}^{N_{e}}P\bigl{(}Y|\text{do}(\hat{C}),E=E_{i}\bigl{)}P\bigl{(}E=E_{i}|\text{do}(\hat{C})\bigl{)}$
$=\sum_{i}^{N_{e}}P\bigl{(}Y|\text{do}(\hat{C}),E=E_{i}\bigl{)}P(E=E_{i})$
$=\sum_{i}^{N_{e}}P\bigl{(}Y|\hat{C},E=E_{i})P(E=E_{i}),$
where $N_{e}$ signifies the count of environment nodes, with $E_{i}$ indicating the $i$-th environmental variable. The environmental enhancement of Cat-GNN is in alignment with the theory of backdoor adjustment, thereby allowing for an effective exploration of possible test environment distributions.
4 Experiments
In this section, we critically assess the CaT-GNN model on a series of research questions (RQs) to establish its efficacy in graph-based fraud detection tasks. The research questions are formulated as follows:
  • RQ1: Does CaT-GNN outperform the current state-of-the-art models for graph-based anomaly detection?
  • RQ2: What is the effectiveness of causal intervention in the aggregation of neighboring information?
  • RQ3: What is the performance with respect to different environmental ratios $r_{e}$?
  • RQ4: Is CaT-GNN equally effective in semi-supervised settings, and how does it perform with limited labeled data?
  • RQ5: Does the causal intervention component lead to a significant decrease in efficiency?
4.1 Experimental Setup
Datasets.
we adopt one private $\mathbf{f}$inacial $\mathbf{f}$raud $\mathbf{s}$emi-supervised $\mathbf{d}$ataset Xiang et al. (2023), termed FFSD, with the partially labeled transaction records. Same with the definition in section 2, if a transaction is reported by a cardholder or identified by financial experts as fraudulent, the label $y_{v}$ will be 1; otherwise, $y_{v}$ will be 0. In addition, we also validate on two other public fraud detection datasets $\mathbf{Yelpchi}$ and $\mathbf{Amazon}$. $\mathbf{Yelpchi}$ Rayana and Akoglu (2015) compiles a collection of hotel and restaurant reviews from Yelp, in which nodes represent reviews. And there are three kinds of relationship edges among these reviews. $\mathbf{Amazon}$: The Amazon graph McAuley and Leskovec (2013) comprises reviews of products in the musical instruments category, in which nodes represent users, and the edges are the corresponding three kinds of relationships among reviews. The statistics of the above three datasets are shown in Table 1.
Table 1: Statistics of the three datasets.
Dataset #Node #Edge #Fraud #benigh
YelpChi 45,954 7,739,912 6,677 39,277
Amazon 11,948 8,808,728 821 11,127
FFSD 1,820,840 31,619,440 33,858 141,861
Baselines.
To verify the effectiveness of our proposed CaT-GNN, we compare it with the following state-of-the-art methods. ❶ Player2Vec. Attributed Heterogeneous Information Network Embedding Framework Zhang et al. (2019). ❷ Semi-GNN. A semi-supervised graph attentive network for financial fraud detection that adopts the attention mechanism to aggregate node embed dings across graphs Wang et al. (2019). ❸ GraphConsis. The GNN-based fraud detectors aim at the inconsistency problem Liu et al. (2020). ❹ GraphSAGE. The inductive graph learning model is based on a fixed sample number of the neighbor nodes Hamilton et al. (2017). ❺ CARE-GNN The camouflage-resistant GNN-based model tackling fraud detection Dou et al. (2020). ❻ PC-GNN. A GNN-based model to address the issue of class imbalance in graph-based fraud detection Liu et al. (2021). ❼ GTAN. A semi-supervised GNN-based model that utilizes a gated temporal attention mechanism to analyze credit card transaction data Xiang et al. (2023). ❽: CaT-GNN (PL). This variant of the CaT-GNN framework selects environment nodes based on a proportion $r_{e}$ and determines mixup weights $a_{i}$ via a learnable linear layer. ❾: CaT-GNN (PI). This version employs a proportional selection of environment nodes and leverages the nodes’ importance scores to inform mixup weights $a_{i}=I_{i}/\sum_{i=1}^{k}I_{i}$. ❿: CaT-GNN (FL). This variant uses a fixed number of environment nodes. Mixup weights are determined by a learnable linear layer. ❿:CaT-GNN (FI). Combining fixed environmental node selection with importance-based weighting for mixup.
Reproducibility
In our experiment, the learning rate $l_{r}$ is set to 0.003, and the batch size batch $N_{batch}$ is established at 256. Moreover, the input dropout ratio $r_{dropout}$ is determined to be 0.2, with the number of attention heads $N_{head}$ set to 4, and the hidden dimension $d$ to 256. We employed the Adam optimizer to train the model over $N_{epoch}=100$ epochs, incorporating an early stopping mechanism to prevent overfitting. In GraphConsis, CARE-GNN, PC-GNN and GTAN, we used the default parameters suggested by the original paper. In Semi-GNN and Player2Vec, We set the learning rate to 0.01. In YelpChi and Amazon datasets, the train, validation, and test ratio are set to be 40%, 20%, and 40% respectively. In the FFSD dataset, we use the first 7 months’ transactions as training data, and the rest as test data. Similar to previous work Liu et al. (2021), we repeat experiments with different random seeds 5 times and we report the average and standard error. Experimental results are statistically significant with $p<0.05$. Cat-GNN and other baselines are all implemented in Pytorch 1.9.0 with Python 3.8. All the experiments are conducted on Ubuntu 18.04.5 LTS server with 1 NVIDIA Tesla V100 GPU, 440 GB RAM.
Metrics.
We selected three representative and extensively utilized metrics: AUC (Area Under the ROC Curve), F1-macro and AP (averaged precision). The first metric AUC is the area under the ROC Curve and as a single numerical value, AUC succinctly summarizes the classifier’s overall performance across all thresholds. The second metric F1-macro is the macro average of F1 score which can be formulated as $F1_{macro}=1/(l\sum_{i=1}^{l}\frac{2\times P_{i}\times R_{i}}{P_{i}+R_{i}})$, and the third metric AP is averaged precision that can be formulated as $AP=\sum_{i=1}^{l}(R_{i}-R_{i-1})P_{i}$, where $P_{i}$ stands for the Precision and $R_{i}$ stands for recall.
Table 2: Performance Comparison (in percent ± standard deviation) on YelpChi, Amazon and FFSD datasets across five runs. The best performances are marked with bold font, and the second-to-best are shown underlined.
Dataset YelpChi Amazon FFSD
Metric AUC F1 AP AUC F1 AP AUC F1 AP
Player2Vec 0.7012$\pm$0.0089 0.4120$\pm$0.0142 0.2477$\pm$0.0161 0.6187$\pm$0.0152 0.2455$\pm$0.0091 0.1301$\pm$0.0117 0.5284$\pm$0.0101 0.2149$\pm$0.0136 0.2067$\pm$0.0155
Semi-GNN 0.5160$\pm$0.0154 0.1023$\pm$0.0216 0.1809$\pm$0.0205 0.7059$\pm$0.0211 0.5486$\pm$0.0105 0.2248$\pm$0.0142 0.5460$\pm$0.0125 0.4393$\pm$0.0152 0.2732$\pm$0.0207
GraphSAGE 0.5414$\pm$0.0029 0.4516$\pm$0.0954 0.1806$\pm$0.0866 0.7590$\pm$0.0053 0.5926$\pm$0.0087 0.6597$\pm$0.0079 0.6534$\pm$0.0095 0.5396$\pm$0.0101 0.3881$\pm$0.0089
GraphConsis 0.7046$\pm$0.0287 0.6023$\pm$0.0195 0.3269$\pm$0.0186 0.8761$\pm$0.0317 0.7725$\pm$0.0319 0.7296$\pm$0.0301 0.6554$\pm$0.0412 0.5436$\pm$0.0376 0.3816$\pm$0.0341
CARE-GNN 0.7745$\pm$0.0281 0.6252$\pm$0.0091 0.4238$\pm$0.0151 0.8998$\pm$0.0925 0.8468$\pm$0.0085 0.8117$\pm$0.0114 0.6589$\pm$0.1078 0.5725$\pm$0.0096 0.4004$\pm$0.0090
PC-GNN 0.7997$\pm$0.0021 0.6429$\pm$0.0205 0.4782$\pm$0.0194 0.9472$\pm$0.0019 0.8798$\pm$0.0084 0.8442$\pm$0.0096 0.6707$\pm$0.0031 0.6051$\pm$0.0230 0.4479$\pm$0.0210
GTAN 0.8675$\pm$0.0036 0.7254$\pm$0.0197 0.6425$\pm$0.0154 0.9580$\pm$0.0014 0.8954$\pm$0.0095 0.8718$\pm$0.0083 0.7496$\pm$0.0041 0.6714$\pm$0.0089 0.5709$\pm$0.0097
Cat-GNN(FI) 0.8721$\pm$0.0044 0.7336$\pm$0.0295 0.6528$\pm$0.0209 0.9643$\pm$0.0026 0.9011$\pm$0.0129 0.8794$\pm$0.0102 0.7643$\pm$0.0078 0.6907$\pm$0.0198 0.5925$\pm$0.0174
Cat-GNN(FL) 0.8910$\pm$0.0026 0.7692$\pm$0.0182 0.6687$\pm$0.0135 0.9705$\pm$0.0016 0.9125$\pm$0.0099 0.8942$\pm$0.0081 0.8023$\pm$0.0067 0.7031$\pm$0.0154 0.6145$\pm$0.0169
Cat-GNN(PI) 0.8895$\pm$0.0041 0.7706$\pm$0.0223 0.6701$\pm$0.0181 0.9669$\pm$0.0021 0.9077$\pm$0.0113 0.8896$\pm$0.0095 0.8145$\pm$0.0061 0.7096$\pm$0.0149 0.6294$\pm$0.0166
Cat-GNN(PL) 0.9035$\pm$0.0035 0.7783$\pm$0.0209 0.6863$\pm$0.0127 0.9706$\pm$0.0015 0.9163$\pm$0.0104 0.8975$\pm$0.0089 0.8281$\pm$0.0054 0.7211$\pm$0.0115 0.6457$\pm$0.0156
4.2 Performance Comparison (RQ1)
In the experiment of credit card fraud detection across three distinct datasets, Cat-GNN showcases superior performance metrics compared to its counterparts. First of all, Cat-GNN achieves the highest AUC in all three datasets, with values of 0.9035, 0.9706, and 0.8281 for YelpChi, Amazon, and FFSD, respectively. This indicates that Cat-GNN consistently outperforms other methods in distinguishing between classes across diverse datasets. Focusing on the F1 Score, which balances the precision $P_{i}$ and recall $R_{i}$, Cat-GNN again tops the charts with scores of 0.7783, 0.9163, and 0.7211 for YelpChi, Amazon, and FFSD. This reflects the model’s robustness in achieving high precision while not compromising on recall, which is essential where both false positives and false negatives carry significant consequences. Finally, Cat-GNN’s superiority extends to the AP metric, with the improvement of at least 6.82%, 2.86%, and 13.10% for YelpChi, Amazon and FFSD respectively.
The comparative performance of Cat-GNN is particularly significant when contrasted with previous methods such as Player2Vec, Semi-GNN, and GraphSAGE. For the Amazon dataset, existing state-of-the-art models, like CARE-GNN, PC-GNN, and GTAN, have already proven effective at capturing the inherent correlations within the data. In this context, the benefits of causal intervention may not be as pronounced, possibly due to the dataset’s simpler local structures and more uniform distribution. However, for the FFSD dataset, our methodology exhibits significant performance improvements. This enhancement is attributed to the complex local structures and the prevalence of unlabeled nodes within the dataset. In such scenarios, causal intervention adeptly learns the inherent attribute connections, thereby boosting the model’s generalization. Additionally, learning mixup weights with a linear layer is more reasonable than weighting with importance scores. Similarly, selecting environment nodes based on proportions is more sensible than choosing a fixed number of environment nodes, and the effect is also slightly better. All in all, This superior performance can be ascribed to the integration of causal theory within the Cat-GNN, enhancing its capacity to comprehend the inherent principles of graph attributes, allowing it to discern complex patterns and interactions that other models are unable to effectively capture.
4.3 Ablation Study (RQ2)
In this section, we evaluate the effectiveness of causal interventions in the aggregation within graph structures. Initially, we explore a variant without any causal intervention, termed N-CaT, which aggregates all neighboring information indiscriminately. Secondly, we introduce D-CaT, a method that omits environment nodes entirely during the aggregation phase, and directly aggregates all neighboring information in the learning process. Finally, our proposed method, CaT, integrates a causal intervention approach, simultaneously considering both causal nodes and environment nodes during aggregation to refine the learning representations.
The results shown in Figure 4 highlight the importance of causal intervention in information aggregation. N-CaT, which lacks causal discernment, performs worse across all datasets compared to CaT because it does not account for causal relationships. D-CaT, which simply deletes environmental factors, shows a significant drop in performance, as the mere deletion of environment nodes prevents the model from fully learning valuable information. Our CaT method consistently outperforms the other variants across all datasets, achieving the highest AUC scores. This superior performance underscores the value of our causal intervention technique, which effectively balances the influence of causal and environment nodes, resulting in a more generalizable model.
Figure 4: The ablation study results on three datasets. Gray bars represent the D-CaT variant, blue bars represent the N-CaT variant, and orange bars represent the CaT-GNN model.
4.4 Parameter Sensitivity Analysis (RQ3, RQ4)
In this section, we study the model parameter sensitivity with respect to the environment nodes ratio and the training ratio. The corresponding results are reported in Figure 5.
Figure 5: Sensitivity analysis with respect to different training ratios (Left) and environment ratios (Right).
As demonstrated in the left of Figure 5, using the YelpChi dataset as an example, the performance of Cat-GNN (measured by AUC as the performance metric) significantly surpasses other competitive models, including PC-GNN and CARE-GNN, across all training ratios, from 10% to 70%. Particularly at lower training ratios (such as 10%), Cat-GNN remains effective for semi-supervised learning and exhibits more robust performance compared to other models.
In our sensitivity analysis of the environmental ratio as demonstrated in the right of Figure 5, we observed that Cat-GNN’s performance on the Amazon dataset is less affected by variations in the training ratio, with AUC fluctuations not exceeding 2%. Conversely, on the FFSD dataset, as the training ratio increases from 5% to 40%, there is a larger fluctuation in Cat-GNN’s performance. This can be attributed to the characteristics of the dataset or the differences in the distribution of labeled data.
4.5 Model Efficiency (RQ5)
In this section, we present a comprehensive analysis of the efficiency of CaT-GNN. Our causal intervention aims to boost performance while maintaining computational efficiency. Table 3 shows that the performance enhancements are achieved without imposing significant additional computational costs. The results indicate that the execution time with causal intervention experienced only a marginal increase. This negligible rise in time is a testament to the algorithm’s ability to retain its computational efficiency while incorporating our advancements. Thus, our algorithm stands as a robust solution that can cater to the needs of high-performance computing while facilitating enhancements that do not compromise on efficiency.
Table 3: Experimental run times with and without causal intervention on three datasets. The experiments were conducted on a Tesla V100 40GB GPU, with the execution times measured in seconds.
Dataset YelpChi Amazon FFSD
No-intervention 126.676 110.518 208.085
Causal-intervention 129.481 (+2.21%) 113.660 (+2.84%) 213.341 (+2.52%)
5 Related Works
Graph Neural Network and its Variants.
Graph neural networks have been widely used in structured data prediction Abadal et al. (2021); Wu et al. (2020); Jiang et al. (2024) by integrating graph structure and attribute. With the development of GNNs, there are several types of GNNs nowadays. 1): Recurrent Graph Neural Networks (RecGNNs) aim to learn node representations with recurrent neural architectures: Scarselli et al. (2008); Gallicchio and Micheli (2010); Li et al. (2015); Dai et al. (2018). 2): Convolutional Graph Neural Networks (ConvGNNs) generalize the operation of convolution from grid data to graph data: Li et al. (2018); Zhuang and Ma (2018); Xu et al. (2018); Chiang et al. (2019). 3): Spatial–Temporal Graph Neural Networks (STGNNs) aim to learn complex hidden patterns from spatial-temporal graphs: Yan et al. (2018); Wu et al. (2019); Guo et al. (2019); Jiang et al. (2023); Li et al. (2022).
Machine-learning based credit card fraud detection.
Recently, numerous efforts have been dedicated to integrating machine learning methodologies into the research of credit card fraud detection. For instance, Maes et al. (2002) successfully applied Bayesian Belief Networks and MLP to the Europay International dataset. Şahin and Duman (2011) utilized decision trees and support vector machines on data from a major national bank. Fu et al. (2016) demonstrates that convolutional neural networks outperform traditional approaches in pattern recognition for higher accuracy. However, their models were limited as they only considered individual transactions or cardholders, missing out on the potential of unlabeled data in real-world transactions Xiang et al. (2023).
GNN-based credit card fraud detection.
More recently, the focus has shifted towards graph-based approaches as the use of graph convolutional networks on datasets with partial labels has been effective for predicting node attributes within citation networks so that many GNN-based fraud detectors have been proposed to detect fraud Wang et al. (2019); Liu et al. (2018). Concretely, Dou et al. (2020) introduces CARE-GNN for fraud detection on relational graphs, while Liu et al. (2021) develops PC-GNN for managing imbalanced learning on graphs. Additionally, Fiore et al. (2019) presents a generative adversarial network to enhance classification capabilities. Cheng et al. (2020) suggested a joint feature learning model, concentrating on spatial and temporal patterns. However, these methods often lack the capability to uncover the causal nature of each specific case and are easily influenced by the surrounding neighbors due to the aggregation mechanism inherent in GNNs. Towards this end, we take the first step to propose a structural causal GNN model, which introduces causal intervention and data augmentation mechanism into the aggregation process of GNN.
6 Conclusion & Future Work
In this work, we introduce the Causal Temporal Graph Neural Network (CaT-GNN), a causal approach in the domain of credit card fraud detection. Our model innovates by integrating causal learning principles to discern and leverage the intricate relationships within transaction data. We validate the effectiveness of CaT-GNN through comprehensive experiments on diverse datasets, where it consistently outperforms existing techniques. Notably, CaT-GNN not only enhances detection accuracy but also maintains computational efficiency, making it viable for large-scale deployment. Future directions will explore extending this methodology to a broader range of fraudulent activities, with the aim of fortifying the integrity of financial systems globally.
Appendix A Future Work
In this paper, we have pioneered the research into integrating causal reasoning with credit card fraud data, systematically constructing a graph neural network solution that addresses out-of-distribution generalization issues caused by local structure inconsistencies in context. Recognizing the potential of causal theory in the development of AI, our future work will delve into more promising forms of data augmentation. For instance, 1): in the field of spatio-temporal data mining, we plan to explore the application of generative diffusion models Guo et al. [2023] for environmental part enhancement; 2) leverage Language Model Learning Liu et al. [2023]; Fang et al. [2024b] for more sophisticated data descriptions and to guide the augmentation of this textual information; 3) graph pruning: Wang et al. [2023b]; Xia et al. [2023]; Wang et al. [2023c, a]; 4) graph sparification: Wu et al. [2023]; Li et al. [2023]; Wang et al. [2024]; Zhang et al. [2024]; Fang et al. [2024a]; 5): graph explainability Fang et al. [2023a, c, 2022]. Furthermore, the method for identifying causal regions can draw on the OAR metric Fang et al. [2023b].

### v2

@@
 Datasets.
-we adopt one private $\mathbf{f}$inacial $\mathbf{f}$raud $\mathbf{s}$emi-supervised $\mathbf{d}$ataset Xiang et al.
-(2023), termed FFSD, with the partially labeled transaction records.
+we adopt one open-source $\mathbf{f}$inacial $\mathbf{f}$raud $\mathbf{s}$emi-supervised $\mathbf{d}$ataset Xiang et al.
+(2023), termed S-FFSD [^2] https://github.com/AI4Risk/antifraud, with the partially labeled transaction records.
 Same with the definition in section 2, if a transaction is reported by a cardholder or identified by financial experts as fraudulent, the label $y_{v}$ will be 1; otherwise, $y_{v}$ will be 0.
@@
 Amazon 11,948 8,808,728 821 11,127
-FFSD 1,820,840 31,619,440 33,858 141,861
+S-FFSD 130,840 3,492,226 2,950 17,553
 Baselines.
@@
 In YelpChi and Amazon datasets, the train, validation, and test ratio are set to be 40%, 20%, and 40% respectively.
-In the FFSD dataset, we use the first 7 months’ transactions as training data, and the rest as test data.
+In the S-FFSD dataset, we use the first 7 months’ transactions as training data, and the rest as test data.
 Similar to previous work Liu et al.
@@
 The second metric F1-macro is the macro average of F1 score which can be formulated as $F1_{macro}=1/(l\sum_{i=1}^{l}\frac{2\times P_{i}\times R_{i}}{P_{i}+R_{i}})$, and the third metric AP is averaged precision that can be formulated as $AP=\sum_{i=1}^{l}(R_{i}-R_{i-1})P_{i}$, where $P_{i}$ stands for the Precision and $R_{i}$ stands for recall.
-Table 2: Performance Comparison (in percent ± standard deviation) on YelpChi, Amazon and FFSD datasets across five runs.
+Table 2: Performance Comparison (in percent ± standard deviation) on YelpChi, Amazon and S-FFSD datasets across five runs.
 The best performances are marked with bold font, and the second-to-best are shown underlined.
-Dataset YelpChi Amazon FFSD
+Dataset YelpChi Amazon S-FFSD
 Metric AUC F1 AP AUC F1 AP AUC F1 AP
@@
 In the experiment of credit card fraud detection across three distinct datasets, Cat-GNN showcases superior performance metrics compared to its counterparts.
-First of all, Cat-GNN achieves the highest AUC in all three datasets, with values of 0.9035, 0.9706, and 0.8281 for YelpChi, Amazon, and FFSD, respectively.
+First of all, Cat-GNN achieves the highest AUC in all three datasets, with values of 0.9035, 0.9706, and 0.8281 for YelpChi, Amazon, and S-FFSD, respectively.
 This indicates that Cat-GNN consistently outperforms other methods in distinguishing between classes across diverse datasets.
-Focusing on the F1 Score, which balances the precision $P_{i}$ and recall $R_{i}$, Cat-GNN again tops the charts with scores of 0.7783, 0.9163, and 0.7211 for YelpChi, Amazon, and FFSD.
+Focusing on the F1 Score, which balances the precision $P_{i}$ and recall $R_{i}$, Cat-GNN again tops the charts with scores of 0.7783, 0.9163, and 0.7211 for YelpChi, Amazon, and S-FFSD.
 This reflects the model’s robustness in achieving high precision while not compromising on recall, which is essential where both false positives and false negatives carry significant consequences.
-Finally, Cat-GNN’s superiority extends to the AP metric, with the improvement of at least 6.82%, 2.86%, and 13.10% for YelpChi, Amazon and FFSD respectively.
+Finally, Cat-GNN’s superiority extends to the AP metric, with the improvement of at least 6.82%, 2.86%, and 13.10% for YelpChi, Amazon and S-FFSD respectively.
 The comparative performance of Cat-GNN is particularly significant when contrasted with previous methods such as Player2Vec, Semi-GNN, and GraphSAGE.
@@
 In this context, the benefits of causal intervention may not be as pronounced, possibly due to the dataset’s simpler local structures and more uniform distribution.
-However, for the FFSD dataset, our methodology exhibits significant performance improvements.
+However, for the S-FFSD dataset, our methodology exhibits significant performance improvements.
 This enhancement is attributed to the complex local structures and the prevalence of unlabeled nodes within the dataset.
@@
 In our sensitivity analysis of the environmental ratio as demonstrated in the right of Figure 5, we observed that Cat-GNN’s performance on the Amazon dataset is less affected by variations in the training ratio, with AUC fluctuations not exceeding 2%.
-Conversely, on the FFSD dataset, as the training ratio increases from 5% to 40%, there is a larger fluctuation in Cat-GNN’s performance.
+Conversely, on the S-FFSD dataset, as the training ratio increases from 5% to 40%, there is a larger fluctuation in Cat-GNN’s performance.
 This can be attributed to the characteristics of the dataset or the differences in the distribution of labeled data.
@@
 Table 3: Experimental run times with and without causal intervention on three datasets. The experiments were conducted on a Tesla V100 40GB GPU, with the execution times measured in seconds.
-Dataset YelpChi Amazon FFSD
+Dataset YelpChi Amazon S-FFSD
 No-intervention 126.676 110.518 208.085
@@
 Future directions will explore extending this methodology to a broader range of fraudulent activities, with the aim of fortifying the integrity of financial systems globally.
-Appendix A Future Work
-In this paper, we have pioneered the research into integrating causal reasoning with credit card fraud data, systematically constructing a graph neural network solution that addresses out-of-distribution generalization issues caused by local structure inconsistencies in context.
-Recognizing the potential of causal theory in the development of AI, our future work will delve into more promising forms of data augmentation.
-For instance, 1): in the field of spatio-temporal data mining, we plan to explore the application of generative diffusion models Guo et al.
-[2023] for environmental part enhancement; 2) leverage Language Model Learning Liu et al.
-[2023]; Fang et al.
-[2024b] for more sophisticated data descriptions and to guide the augmentation of this textual information; 3) graph pruning: Wang et al.
-[2023b]; Xia et al.
-[2023]; Wang et al.
-[2023c, a]; 4) graph sparification: Wu et al.
-[2023]; Li et al.
-[2023]; Wang et al.
-[2024]; Zhang et al.
-[2024]; Fang et al.
-[2024a]; 5): graph explainability Fang et al.
-[2023a, c, 2022].
-Furthermore, the method for identifying causal regions can draw on the OAR metric Fang et al.
-[2023b].